Steam enters a turbine from a height of 6 m, at 50m/s speed at 2MPa pressure and 350°C, and after 8.5kW heat loss to the environment, it exits the turbine at 0.1Mpa pressure, with a dryness degree of 0.92, from a height of A meter at a speed of 200m/s. (A indicates a single number height)

Calculate the power of the turbine if the mass of steam entering the turbine is 1.5kg/s.


1
Expert's answer
2021-12-27T07:54:41-0500

"p_0=\rho g(h_1-h_2)+\frac{\rho (v_1^2-v_2^2)}2+(p_1-p_2)=1.56~MPa,"

"A=\eta E-\frac{\dot m}M RT=7.4~kJ."